OpenStudy (anonymous):

There are 9 Republicans and 5 Democrats on a committee. A sub committee of four is selected at random. What is the probability that all are Republicans, given that at least two are Republicans?

OpenStudy (anonymous):

yeah,, but the teachers love to give out these kind of question..

OpenStudy (anonymous):

Poorly worded question.

OpenStudy (anonymous):

First find the probability that two are republicans.

OpenStudy (♪chibiterasu):

Well we have 9 republicans, and there is a probability of there being four of them. All we have to do is put 4 on top of the total of republicans and Democratsand we have our solution: 9 + 5 = 14 4/14 Now we simplify: 4/14 2/7 There is a 2/7 chance of there being all republicans chosen.

OpenStudy (anonymous):

Then find the probability that 4 or republicans.

OpenStudy (♪chibiterasu):

Ah, I didn't see that last part.

OpenStudy (♪chibiterasu):

Well we just take away two from the total: 4/12 2/6 1/3 There is a 1/3 chance of there being all republicans.

OpenStudy (anonymous):

Okay the probability that at least 2 are republicans is a bit more complicated.

OpenStudy (anonymous):

Its the probability that 2, 3, 4 republicans are in it.

OpenStudy (anonymous):

So for exactly 2 it is:\[ \frac{9}{14}\times\frac{8}{13}\times\frac{5}{12}\times\frac{4}{11}=\frac{60}{1001} \]For exactly 3 it is: \[ \frac{9}{14}\times\frac{8}{13}\times\frac{7}{12}\times\frac{5}{11}=\frac{5}{143} \]For exactly 4 it is: \[ \frac{9}{14}\times\frac{8}{13}\times\frac{7}{12}\times\frac{6}{11}=\frac{18}{143} \] Add all of these up.\[ \frac{60}{1001}+\frac{5}{143}+\frac{18}{143}=\frac{17}{77} \] The conditional probability is just the probability of exactly 4 AND at least 2 divided by the probability of at least two. At least 2 and exactly 4 is not different than exactly 4 because exactly 4 implies at least 2. So we have: \[ \frac{18}{143}\div\frac{17}{77}=\frac{126}{221}\approx 57\% \]

OpenStudy (anonymous):

A tedious question indeed.